Evaluate det(−A) and det(−A T ). Compare det(−A) and det(−A T ) for:

(2.1) A = -4 2

3 -3


(2.2) A = 3 1 -2

-5 3 -6

-1 0 -4


1
Expert's answer
2021-05-31T16:26:56-0400

The determinant of the transpose of a square matrix is equal to the determinant of the matrix, that is,


"\det(A^T)=\det (A)"

2.1


"A=\begin{bmatrix}
 -4 & 2 \\
 3 & -3
\end{bmatrix}"

"-A=\begin{bmatrix}
 4 & -2 \\
 -3 & 3
\end{bmatrix}"

"\det (-A)=\begin{vmatrix}
 4 & -2 \\
 -3 & 3
\end{vmatrix}=4(3)-(-2)(-3)=18"

"A^T=\begin{bmatrix}
 -4 & 2 \\
 3 & -3
\end{bmatrix}^T=\begin{bmatrix}
 -4 & 3 \\
 2 & -3
\end{bmatrix}"

"-A^T=\begin{bmatrix}
 4 & -3 \\
 -2 & 3
\end{bmatrix}"


"\det (-A^T)=\begin{vmatrix}
 4 & -3 \\
 -2 & 3
\end{vmatrix}=4(3)-(-2)(-3)=18"


"\det (-A)=\det (-A^T)"

2.2



"A=\begin{bmatrix}
 3 & 1 & -2 \\
 -5 & 3 & -6 \\
-1 & 0 & -4
\end{bmatrix}"

"-A=\begin{bmatrix}
 -3 & -1 & 2 \\
 5 & -3 & 6 \\
1 & 0 & 4
\end{bmatrix}"

"\det (-A)=\begin{vmatrix}
 -3 & -1 & 2 \\
 5 & -3 & 6 \\
1 & 0 & 4
\end{vmatrix}"

"=1\begin{vmatrix}
 -1& 2 \\
 -3 & 6
\end{vmatrix}-0\begin{vmatrix}
 -3 & 2 \\
 5 & 6
\end{vmatrix}+4\begin{vmatrix}
 -3 & -1 \\
 5 & -3
\end{vmatrix}"

"=-6+6+4(9+5)=56"

"A^T=\begin{bmatrix}
 3 & 1 & -2 \\
 -5 & 3 & -6 \\
-1 & 0 & -4
\end{bmatrix}^T=\begin{bmatrix}
 3 &- 5 & -1 \\
 1 & 3 & 0 \\
-2 & -6 & -4
\end{bmatrix}"

"-A^T=\begin{bmatrix}
 -3 & 5 & 1 \\
 -1 & -3 & 0 \\
2 & 6 & 4
\end{bmatrix}"


"\det (-A^T)=\begin{vmatrix}
 -3 & 5 & 1 \\
 -1 & -3 & 0 \\
2 & 6 & 4
\end{vmatrix}"

"=1\begin{vmatrix}
 -1& -3 \\
 2 & 6
\end{vmatrix}-0\begin{vmatrix}
 -3 & 5 \\
 2 & 6
\end{vmatrix}+4\begin{vmatrix}
 -3 &5 \\
 -1 & -3
\end{vmatrix}"


"=-6+6-9+4(9+5)=56"


"\det (-A)=\det (-A^T)"